# DriveVLA-W0: World Models Amplify Data Scaling Law in Autonomous Driving

### v1

@@ 4.3 Comparison with State-of-the-art Methods
As presented in Table 1 and Table 2, DriveVLA-W0 establishes a new state-of-the-art on the NAVSIM benchmark by surpassing top-performing methods across different architectural paradigms, including the BEV-based WoTE (Li et al., 2025a) and the VLA-based AutoVLA (Zhou et al., 2025c). Notably, our model achieves this top performance using only a single front-view camera, surpassing competitors that rely on richer sensor suites combining multi-view cameras and LiDAR. This superior performance is attributed to the powerful, dense supervision from our world modeling objective, which enables the model to learn a more effective feature representation.
@@ 4.4 World Models Amplify Data Scaling Law
Figure 4: World modeling unlocks generalization with data scaling. Our world model turns pretraining from a detriment for sparse-supervision baselines (red arrows) into a benefit for our VLA-W0s (green arrows), enabling positive knowledge transfer across datasets with similar visuals (b) but different action distributions (a). Figure (a) is from Dauner et al. (2024).
@@ 4.5 Ablation Study
Pretrain
Finetune
NC $\uparrow$ DAC $\uparrow$ PDMS $\uparrow$
@@
Pretrain
Finetune
NC $\uparrow$ DAC $\uparrow$ PDMS $\uparrow$
@@
Ablation on Latency We validate the efficiency of MoE architecture by measuring inference latency on an H200 GPU. Compared to the baseline DriveVLA-W0 (117.8ms latency, 85.6 PDMS), the addition of our query-based MoE expert significantly cuts the latency to 74.3ms (just 63.1% of the original) while simultaneously boosting performance to 88.4 PDMS. More analysis is shown in §B.1.
@@ B.2 Positive Correlation Between Generative Fidelity and Planning Performance
Sequence Design
FID
$\downarrow$
PDMS
$\uparrow$
@@ Appendix C Visualization
C.1 Case Analysis
In this section, we provide a qualitative case analysis to offer deeper insights into our model’s behavior and validate our key findings. Our analysis is twofold. First, as shown in Figure 6, we compare our VLA-W0 against baseline models (VLA baseline and TransFuser) in a challenging corner case to visually demonstrate the benefits of world modeling for navigating complex scenarios. Second, we compare the trajectories generated by different action experts built upon the same VLA backbone. The subsequent visualizations in Figure 7,8,9 reveal that the AR expert generates markedly more stable trajectories than the flow-matching approach, particularly in terms of inter-frame consistency and avoiding aggressive maneuvers.
@@ C.1 Case Analysis
C.2 Image Generation by World Model
Our world model demonstrates both high visual fidelity and strong action consistency. As shown in Figure 10, it generates realistic and plausible images across diverse and challenging scenarios. More importantly, these visual predictions are tightly coupled with the model’s planning process. As illustrated in Figure 11, any systematic errors in the predicted trajectory (e.g., a rightward deviation) are consistently mirrored in the generated images (a rightward shift). This correspondence indicates that the model’s planned actions are closely grounded in its internal visual predictions, reflecting a coherent and predictive understanding of the environment.
Figure 10: Future image generation. Our world model demonstrates strong generative fidelity, producing visually realistic and contextually plausible futures across diverse and challenging scenarios, including complex intersections, dense traffic, and nighttime conditions.
Figure 11: Alignment between predicted images and actions. The model demonstrates a tight coupling between its visual predictions and action predictions.
@@ Appendix E More Related Work
World Models Research on world models has mainly followed two distinct philosophies: using them as powerful data synthesizers for generating realistic scenarios, or as an auxiliary objective for representation learning to aid downstream tasks. The first and more common approach focuses on generation. Many works in autonomous driving aim to synthesize high-quality driving data. For instance, some models generate future images directly. GAIA-1 (Hu et al., 2023) uses video, text, and actions to create realistic scenarios; MILE (Hu et al., 2022a) leverages 3D geometry as an inductive bias; and DrivingGPT (Chen et al., 2024) generates both future images and corresponding actions. Other works like Copilot4D (Zhang et al., 2023) predict future discrete visual tokens. A related line of research focuses on generating future spatial representations like occupancy grids, as seen in OccWorld (Zheng et al., 2023) and DriveWorld (Min et al., 2024). In contrast, a second, innovative line of research leverages world modeling as a self-supervised objective to enhance representation learning. This is particularly effective in robotics, where UniVLA (Wang et al., 2025) and WorldVLA (Cen et al., 2025) use next-token prediction in an Autoregressive VLA to learn robust policies. In the context of autonomous driving, LAW (Li et al., 2024a) pioneers this approach by using a latent world model to assist end-to-end feature learning. However, LAW is designed to predict future states in a latent space. Our work is distinct in that we supervise the model by predicting future images, providing a richer and more direct learning signal for the VLA backbone.

### v2

@@ 4.3 Comparison with State-of-the-art Methods
As presented in Table 1 and Table 2, DriveVLA-W0 establishes a new state-of-the-art on the NAVSIM benchmark by surpassing top-performing methods across different architectural paradigms, including the BEV-based WoTE (wote) and the VLA-based AutoVLA (zhou2025autovla). Notably, our model achieves this top performance using only a single front-view camera, surpassing competitors that rely on richer sensor suites combining multi-view cameras and LiDAR. This superior performance is attributed to the powerful, dense supervision from our world modeling objective, which enables the model to learn a more effective feature representation.
@@ 4.4 World Models Amplify Data Scaling Law
Figure 4: World modeling unlocks generalization with data scaling. Our world model turns pretraining from a detriment for sparse-supervision baselines (red arrows) into a benefit for our VLA-W0s (green arrows), enabling positive knowledge transfer across datasets with similar visuals (b) but different action distributions (a). Figure (a) is from navsim.
@@ 4.5 Ablation Study
Pretrain (NuPlan) Finetune (NAVSIM) NC $\uparrow$ DAC $\uparrow$ PDMS $\uparrow$
@@
Pretrain (NuPlan) Finetune (NAVSIM) NC $\uparrow$ DAC $\uparrow$ PDMS $\uparrow$
@@
Ablation on Latency We validate the efficiency of MoE architecture by measuring inference latency on an H200 GPU. Compared to the baseline DriveVLA-W0 (117.8ms latency, 85.6 PDMS), the addition of our query-based MoE expert significantly cuts the latency to 74.3ms (just 63.1% of the original) while simultaneously boosting performance to 88.4 PDMS. More analysis is shown in AppendixB.1.
4.6 Visualization
Beyond achieving superior planning performance, DriveVLA-W0 exhibits strong capabilities in action-conditioned future generation, highlighting its potential as a reactive simulator. As illustrated in Figure 5, we explicitly condition the model on a counterfactual “decelerate” trajectory. The resulting frames show the surrounding scene flowing past the vehicle at a noticeably reduced rate compared to the ground truth. This precise alignment between control signals and visual dynamics confirms the model’s grounded physical understanding. For further qualitative analysis, including trajectory visualization and case comparisons (Figures 7, 8, 9, 10), future image generation (Figure 13), and counterfactual reasoning (Figure 14), please refer to Appendix C.
Figure 5: Action-Conditioned Simulation. We demonstrate the utility of our world model as a simulator capable of generating diverse futures based on control signals. In the recorded ground truth, the ego vehicle maintains speed and drives forward. By conditioning the model on a counterfactual "decelerate" action, the world model generates a distinct visual outcome where the vehicle slows down.
@@ B.2 Positive Correlation Between Generative Fidelity and Planning Performance
Sequence Design FID $\downarrow$ PDMS $\uparrow$
@@ Appendix C Visualization
C.1 Planning Trajectory
In this section, we provide a qualitative case analysis to offer deeper insights into our model’s behavior and validate our key findings. Our analysis is twofold. First, as shown in Figure 7, we compare our DriveVLA-W0 against baseline models (VLA baseline and TransFuser) in a challenging corner case to visually demonstrate the benefits of world modeling for navigating complex scenarios. Second, we compare the trajectories generated by different action experts built upon the same VLA backbone. The subsequent visualizations in Figure 8,9,10 reveal that the AR expert generates markedly more stable trajectories than the flow-matching approach, particularly in terms of inter-frame consistency and avoiding aggressive maneuvers.
@@ C.1 Planning Trajectory
C.2 Failure Case
To provide a balanced view of our model’s performance and identify directions for future improvement, we analyze representative failure cases encountered during evaluation. We categorize these failures into two primary types: instruction ambiguity and dynamic prediction errors. As illustrated in Figure 11, coarse-grained navigation commands can lead to indecision in complex road topologies (e.g., Y-junctions). Additionally, Figure 12 highlights the challenge of predicting fine-grained dynamic objects (e.g., oncoming vehicles) in high-interaction scenarios.
Figure 11: Failure case analysis: ambiguity in driving commands. We illustrate a failure mode stemming from the ambiguity of coarse-grained navigation instructions. In this Y-shaped intersection, the model receives a generic “go straight” command. Due to the splitting road topology, this instruction is ambiguous, as it maps clearly to neither the left nor the right branch. Consequently, the model fails to commit to a valid lane, hesitantly driving straight into the fork area (as shown by the predicted red trajectory). This highlights a limitation in the current NAVSIM benchmark, where the discrete command set (restricted to simple Turn Left/Straight/Right) lacks the granularity to resolve ambiguities in complex geometries. Note: To faithfully represent the visual information available to the model, all displayed images are reconstructions from the MoVQGAN tokens, rather than raw RGB frames.
Figure 12: Failure case analysis: dynamic object prediction. While our world model generalizes well to most dynamic scenes, it faces challenges in predicting multiple objects at complex intersections. In this specific case, the world model fails to anticipate the emergence of oncoming vehicles in the future frames (as reflected in the Cond. on GT Traj). Consequently, the planner, unaware of the potential conflict, incorrectly executes a left turn. This highlights the modeling of fine-grained dynamic objects as a valuable direction for future research. Note: To faithfully represent the visual information available to the model, all displayed images are reconstructions from the MoVQGAN tokens, rather than raw RGB frames.
C.3 Future Image Generation
Our world model demonstrates both high visual fidelity and strong action consistency. As shown in Figure 13, it generates realistic and plausible images across diverse and challenging scenarios. More importantly, these visual predictions are tightly coupled with the model’s planning process.
Figure 13: Future image generation. Our world model demonstrates strong generative fidelity, producing visually realistic and contextually plausible futures across diverse and challenging scenarios, including complex intersections and dense traffic.
C.4 Counterfactual Reasoning
This section investigates the model’s capability for counterfactual reasoning, synthesizing alternative futures based on hypothetical actions. As shown in Figure 14, conditioning the model on a "turn right" action generates realistic off-road imagery despite the ground truth being straight. This confirms that the model captures the underlying scene geometry rather than merely memorizing training trajectories.
Figure 14: Counterfactual reasoning. By conditioning on a counterfactual "turn right" action at the roadside, our world model generates realistic video sequences where the vehicle deviates from the ground truth trajectory and drives off-road. This validates that the model has learned the underlying scene geometry rather than simply memorizing the training data.
@@ Appendix E More Related Work
Scaling Laws in Deep Learning. kaplan2020scaling were the first to systematically demonstrate that pretraining loss scales as a power law with model size, dataset size, and computational cost. Chinchilla (hoffmann2022training) then showed many LMs were undertrained and derived a compute-optimal prescription that scales model size and tokens proportionally. In computer vision, zhai2022scaling charted ViT scaling law with stable training recipes, and ViT-22B (dehghani2023scaling) scaled ViTs to 22B parameters, verifying predictable multi-task improvements. lindata conducted a large-scale study of imitation-learning data scaling in robotics, and found near–power-law gains from increasing environmental and object diversity with improved zero-shot generalization. In autonomous driving, STR (sun2023large) shows large trajectory models scale steadily in both prediction and planning, and baniodeh2025scaling reported power-law improvements for joint motion forecasting and planning with large driving datasets. For end-to-end driving, naumann2025data observe roughly log-linear gains in both open- and closed-loop metrics as training data scale increases. zheng2024preliminary also observe power-law improvements from data scaling in a large-scale imitation learning study. However, zheng2024preliminary exclusively analyzes scaling behavior within the conventional paradigm of sparse action supervision. Our work, in contrast, provides the first study of how these scaling laws are reshaped by the self-supervised signals provided by world modeling.